OpenStudy (anonymous):

suppose you draw a card randomly from a normal deck of cards. What is the probability you will draw a black 4?

OpenStudy (anonymous):

@twvogels

OpenStudy (anonymous):

assuming 2 black 4's in a 52 deck card the probability is 2/52 or 1/26

OpenStudy (anonymous):

The total number of cards in a deck is 52. So there are 52 total possible outcomes. There are two Black 4s in a deck. So there are 2 'correct' outcomes. To find the probability we do\[\frac{correct}{total}=\frac{2}{52}=\frac{1}{26}\] Which is also about 0.0385 or 3.85%.

OpenStudy (anonymous):

Thanks!